# QuickVideo: Real-Time Long Video Understanding with System Algorithm Co-Design

### v1

@@ 4.2 QuickPrefill Results
In this section, we evaluate the performance of QuickPrefill across four long video understanding benchmarks whose video length ranges from minutes to hours: VideoMME [14], LongVideoBench [40], LVBench [34], and MLVU [46]. All generations are conducted using greedy sampling, and we report results using the lmms-eval framework [44]. All experiments are performed using the Qwen2.5-VL-7B-Instruct model [4] on a single A100 (40GB) GPU with 8 replicas.
@@ 64 Frames
Figure 4: Ablation study on the group size and retention ratio. Data from Table 2.
@@
We evaluate the impact of various KV cache pruning strategies on model accuracy, as summarized in Table 1. Specifically, we compare several pruning techniques against a baseline with no pruning applied ($\rho=1$). We fix the KV cache retention ratio $\rho$ to 0.5 and set the group size to 16 frames.
Among the evaluated strategies, the Key Norms (small) method achieves the best trade-off between efficiency and accuracy, retaining over 95% of the model’s original performance while reducing KV cache size and computation by half. In the 1024 frames setting, it can even retain more than 98% of the model’s original performance Notably, this method outperforms alternatives that select tokens based on query attention scores. While prior work [11] demonstrated that negative L2 norms of keys correlate strongly with attention scores in text-only LLMs, our results extend this finding to the VideoLLM prefilling setting. This suggests that key norm-based pruning is not only effective during decoding but also applicable in the context of prefilling, underscoring its generalizability and practical value for long-context video understanding.
We further conduct ablation studies to assess the effects of group size and the retention ratio $\rho$ (see Appendix E). As shown in Table 2 and Figure 4, varying the group size has minimal impact on model performance, while increasing the retention ratio $\rho$ consistently improves accuracy, eventually matching the performance of the non-pruned baseline. Smaller group sizes yield lower activation memory, and lower retention ratios result in reduced KV cache memory. These findings offer practical guidance for balancing memory efficiency and model accuracy based on user constraints.
@@ 4.3 Latency in end-to-end QuickVideo Inference
Figure 5: Latency from video loading, prefill and LLM decoding in an end-to-end inference setting. We compare a baseline implementation of Qwen2.5-VL [4], Qwen2.5-VL implemented with QuickPrefill and QuickCodec, and lastly, our block overlapped design.
We implement both QuickCodec and QuickPrefill into a Qwen2.5-VL-7B-Instruct [4] inference pipeline. We implement our two versions: loading the entire video input using QuickCodec before prefilling with QuickPrefill, as well as our group overlapped variant. We benchmark latency from the video loading, prefill and LLM decoding stages in our end-to-end inference pipeline. For QuickPrefill, we use the Key Norms (small) as the KV cache pruning method and set $\rho$=0.2, and the number of frames in each group to be 32. We use a 30 minute video (sampled at 1 FPS), as the baseline implementation runs out of memory when using longer video inputs. We use an A100 80GB SXM4 GPU and an AMD EPYC 7513 32-Core CPU. We allocate 16 cores for video processing. For our overlapped implementation, we use 64 video segments ($s$) for our parallelized loading. Our timings are averaged over 10 runs.
Figure 5 shows the latency from video loading, prefill and LLM decoding of all 3 implementations. After QuickCodec reduces the video loading time, we can almost completely overlap video loading and prefill. In the overlapped pipeline, video processing, prefill and LLM decoding completes in only 20.0 seconds, a 49.7 second speedup over the baseline’s 69.7 seconds. Our overlapped implementation has a small startup latency, due to metadata parsing and decoding the frames for the first prefill block. This amounts to 2.8 seconds of video loading that cannot be overlapped with prefill.
@@ Appendix A Parallelized Interval Algorithm
8:         if $|\mathcal{K}_{j-1}-pts_{estimate}|<|\mathcal{K}_{j}-pts_{estimate}$ then
@@
After finding the locations of keyframes on line 2, Keyframe Intervals computes $c$ intervals as follows: We calculate the length of $\frac{1}{c}$ of the stream, in pts units (line 4). On lines 5-10, we search for the keyframes closest to being $\frac{i}{c}th$ through the video, given by $pts_{estimate}$. FindInsertionIndex uses binary search to find where in the list of keyframes $pts_{estimate}$ would be inserted. After finding the insertion point $j$, the algorithm checks whether the keyframe before or after $j$ is closer to $pts_{estimate}$. The closest keyframe location is added to $\mathcal{J}$, the list of intervals. $\mathcal{J}[0]=pts_{min}$ and $\mathcal{J}[c-1]=pts_{max}$, to ensure that the intervals span the video. $\mathcal{J}[1:c-2]$ are keyframe-aligned and equally spaced. Therefore, $\mathcal{J}$, a list containing $c+1$ values, can be interpreted as $c$ intervals: $\mathcal{J}^{\prime}=\{(\mathcal{J}[i],\mathcal{J}[i+1])\;|\;i\in 0,1,\dots,\;c\}$.

### v2

@@ 4.2 QuickPrefill Results
We evaluate QuickPrefill on four long video understanding benchmarks, with videos ranging from minutes to hours: VideoMME [14], LongVideoBench [40], LVBench [34], and MLVU [46]. All generations use greedy sampling, and results are reported via the lmms-eval framework [44]. Experiments are run on the Qwen2.5-VL-7B-Instruct model [4] using a single A100 (40GB) GPU with 8 replicas.
@@ 64 Frames
Figure 4: Ablation study on group size and retention ratio. Data from Table 2.
@@
We evaluate the impact of various KV cache pruning strategies on model accuracy, as summarized in Table 1. We compare several pruning techniques against a no-pruning baseline ($\rho=1$), fixing the retention ratio $\rho$ at 0.5 and the group size at 16 frames. The Key Norms (small) method achieves the best balance between efficiency and accuracy, retaining over 95% of the model’s original performance while halving the KV cache size and computation. In the 1024-frame setting, it retains over 98% of the original performance. Notably, this method outperforms query-attention-based token selection strategies. While prior work [11] has shown that negative L2 norms of keys correlate strongly with attention scores in text-only LLMs, our results extend this finding to VideoLLM prefilling, highlighting the generalizability and practical utility of key norm-based pruning.
We also conduct ablation studies on group size and retention ratio $\rho$ (see Appendix E). As shown in Table 2 and Figure 4, group size has minimal impact on model performance, while increasing $\rho$ consistently improves accuracy, approaching the no-pruning baseline. Smaller group sizes reduce activation memory, while lower $\rho$ values reduce KV cache memory. These findings provide practical guidance for balancing memory efficiency and model accuracy based on system constraints.
@@ 4.3 Latency in End-to-End QuickVideo Inference
Figure 5: Latency breakdown for video loading, prefill, and LLM decoding in end-to-end inference. We compare a baseline Qwen2.5-VL [4] implementation, the same model with QuickPrefill and QuickCodec, and our block-overlapped design.
We integrate QuickCodec and QuickPrefill into a Qwen2.5-VL-7B-Instruct [4] inference pipeline. We evaluate two configurations: (1) loading the entire video with QuickCodec followed by QuickPrefill, and (2) our group-overlapped design. Latency for video loading, prefill, and LLM decoding is benchmarked in an end-to-end pipeline. For QuickPrefill, we use the Key Norms (small) pruning method with $\rho=0.2$ and set the group size to 32 frames. We use a 30-minute video (sampled at 1 FPS) as the baseline implementation runs out of memory with longer videos. Experiments are conducted on an A100 80GB SXM4 GPU with an AMD EPYC 7513 32-core CPU, allocating 16 cores for video processing. For the overlapped implementation, we use 64 intervals ($s$ in section 3.3) for parallelized loading. All timings are averaged over 10 runs. Figure 5 presents latency breakdowns for all three implementations. After applying QuickCodec, we significantly reduce video loading time. By overlapping video loading and prefill, we achieve near-complete overlap of the two stages. The block-overlapped design completes video processing, prefill, and LLM decoding in 20.0 seconds, a 49.7-second speedup over the baseline’s 69.7 seconds. Our overlapped pipeline introduces a small startup latency—2.8 seconds—for metadata parsing and decoding frames for the first prefill block, which cannot be overlapped.
@@ Appendix A Parallelized Interval Algorithm
8:         if $|\mathcal{K}_{j-1}-pts_{estimate}|<|\mathcal{K}_{j}-pts_{estimate}|$ then
@@
After finding the locations of keyframes on line 2, Keyframe Intervals computes $c$ intervals as follows: We calculate the length of $\frac{1}{c}$ of the stream, in pts units (line 4). On lines 5-10, we search for the keyframes closest to being $\frac{i}{c}th$ through the video, given by $pts_{estimate}$. FindInsertionIndex uses binary search to find where in the list of keyframes $pts_{estimate}$ would be inserted. After finding the insertion point $j$, the algorithm checks whether the keyframe before or after $j$ is closer to $pts_{estimate}$. The closest keyframe location is added to $\mathcal{J}$, the list of intervals. $\mathcal{J}[0]=pts_{min}$ and $\mathcal{J}[c-1]=pts_{max}$, to ensure that the intervals span the video. $\mathcal{J}[1],\mathcal{J}[2],\dots\mathcal{J}[{c-2}]$ are keyframe-aligned and equally spaced. Therefore, $\mathcal{J}$, a list containing $c+1$ values, can be interpreted as $c$ intervals: $\mathcal{J}^{\prime}=\{(\mathcal{J}[i],\mathcal{J}[{i+1}])\;|\;i\in 0,1,\dots,\;c\}$.